OpenStudy (anonymous):

How to find the x and y intercept of f(x)=400-4/5x

hero (hero):

To find the x intercept, set y = 0, then solve for x: Remember y = f(x) so you have y = 400 - (4/5)x Set y = 0 0 = 400 - (4/5)x Add (4/5)x to both sides: (4/5)x = 400 Multiply both sides by (5/4) x = (5/4)(400) x = (5)(100) x = 500

hero (hero):

Using the same logic, you'll be able to find the y-intercept.

hero (hero):

btw, @dev165, try not to post the same question twice.